# Search for ultralight dark matter in the SuperMAG high-fidelity dataset

### v1

@@ I Introduction
Discerning the fundamental nature of dark matter (DM) remains one of the most prominent open problems in modern physics. A number of potential DM candidates have been proposed, but one class which has recently received significant attention is ultralight bosonic DM [1, 2]. This class includes DM candidates with masses $m_{\mathrm{DM}}\lesssim 1\,\textrm{eV}$, [^1] We work in natural units where $\hbar=c=1$. which have correspondingly large number densities and so behave as classical fields [3, 4]. Two such candidates which have garnered particular interest are dark-photon dark matter (DPDM) [5, 6, 7, 8] and axionlike-particle dark matter (henceforth simply axion DM) [9, 10, 11, 12, 13]. Both of these candidates may exhibit couplings with electromagnetism [5, 14], and a number of experimental [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33], astrophysical [34, 35, 36, 37, 38, 39, 40, 41, 42], and cosmological probes [43, 44, 45, 46, 47, 48] have been utilized to constrain their parameter spaces.
Few experiments searching for these ultralight DM candidates are sensitive to the very low mass parameter space $m_{\mathrm{DM}}\lesssim 10^{-12}\,\textrm{eV}$. This is because the signal in these experiments typically scales with the size of the experiment $L$, when $\lambda_{\mathrm{DM}}\gg L$ (where $\lambda_{\mathrm{DM}}=2\pi/m_{\mathrm{DM}}$ is the Compton wavelength of the DM). One notable example is a series of experimental searches which utilize Earth as a transducer for ultralight DM [49, 50, 51, 52, 53]. In such contexts, the effective size of the experiment is the radius of Earth $R$, which leads to a large enhancement of the signal.
@@
In this work, we extend the analyses of Refs. [50, 51] to the SuperMAG “high-fidelity” dataset, which exhibits a shorter one-second time resolution, albeit with fewer, $\mathcal{O}(200)$, total stations. This will enable sensitivity to Compton frequencies up to $f_{\mathrm{DM}}\leq 1\,\textrm{Hz}$. As ambient geomagnetic-field noise is typically reduced at higher frequencies, this DM search strategy should be even more powerful at these higher frequencies. In fact, the SNIPE Hunt collaboration recently conducted an independent search for this same DM signal at even higher frequencies, $0.5\,\textrm{Hz}\leq f_{\mathrm{DM}}\leq 5\,\textrm{Hz}$, and was able to achieve competitive results with data from only three magnetically quiet locations [52, 53]. The search in this work leverages the benefits of both low noise at higher frequencies and the large amounts of data that the high-fidelity dataset has to offer. In this work, we search for ultralight DM in the frequency range $10^{-3}\,\textrm{Hz}\leq f_{\mathrm{DM}}\leq 0.98\,\textrm{Hz}$; which complements the existing searches from the SuperMAG low-fidelity dataset and the SNIPE Hunt collaboration.
This work is organized as follows. In Sec. II, we briefly review the magnetic-field signal of ultralight DM that we search for. In Sec. III, we introduce the SuperMAG high-fidelity dataset and review the data-processing techniques which have been applied to it. In Sec. IV, we describe the details of our analysis procedure. The analysis in this work largely follows the same procedure as in Refs. [50, 51], but with important modifications that we highlight in this section. In Sec. V, we present and discuss the resulting constraints on ultralight DM from our search. In Appendix A, we review the properties of vector spherical harmonics (VSH). In Appendix B, we validate our choice of analysis parameters by examining the statistics of the noise in our data. All the code used in this work is publicly available on Github [57].
@@ II Signal
Here $\Omega=(\theta,\phi)$ denotes the geographic coordinates on Earth’s surface, oriented so that $\theta=0$ corresponds to the Geographic North Pole and $\theta=\pi$ to the Geographic South Pole. (Note that while $\phi$ coincides with the usual definition of longitude, $\theta$ is not the usual latitude.) Additionally, $R$ is the radius of Earth, $A^{\prime}_{m}$ are complex amplitudes corresponding to the polarization modes of the dark photon, [^4] In our conventions, these amplitudes relate to the Cartesian components of $\bm{A}^{\prime}$ via $A^{\prime}_{x}=-\frac{1}{\sqrt{2}}\left(A^{\prime}_{+}-A^{\prime}_{-}\right),\quad A^{\prime}_{y}=-\frac{i}{\sqrt{2}}\left(A^{\prime}_{+}+A^{\prime}_{-}\right),\quad A^{\prime}_{z}=A^{\prime}_{0}.$ and $\bm{\Phi}_{\ell m}$ are VSH (see Appendix A). The presence of $f_{d}=1/(\textrm{sidereal day})$ in the time-dependence of the signal arises due to the rotation of Earth. Here the dark-photon amplitudes $A^{\prime}_{m}$ are defined in the inertial celestial frame, while the signal in Eq. (1) is measured in the rotating Earth-fixed frame. Translating between these frames induces a polarization-dependent shift to the frequency of the signal. (In frequency space, this will introduce sidebands at $f=f_{A^{\prime}}\pm f_{d}$ to the signal.)
@@
Since Earth’s field slowly drifts over time, the Gauss coefficients $g_{\ell m}$ and $h_{\ell m}$ are time-dependent. The IGRF-13 model provides values for these coefficients at five-year intervals from 1900 to 2020, along with their current time derivative. These values can be interpolated to find the values of the Gauss coefficients at any time between the beginning of 1900 and the beginning of 2020, and the time derivatives can be used to extrapolate their values up to the end of 2024.
@@
Before proceeding, we note a few distinctions between the signals in Eqs. (1) and (5). Firstly, the DPDM signal is purely dipolar (only $\ell=1$ modes contribute). Because $\bm{B}_{\oplus}$ is approximately dipolar, the expansion in Eq. (3) is dominated by the $C_{10}$ term, but other terms contribute at the $\mathcal{O}(10\%)$ level. Likewise, the axion signal in Eq. (5) is predominantly dipolar, but receives small contributions from higher $\ell$ modes. Secondly, the axion signal in Eq. (5) does not feature any dependence on $f_{d}$. This is because the angular dependence of the axion signal is set by $\bm{B}_{\oplus}$ which co-rotates with Earth, as opposed to the DPDM polarization which does not. This means that, in frequency space, the axion signal will only appear at $f=f_{a}$, whereas the DPDM signal will appear at $f=f_{A^{\prime}},f_{A^{\prime}}\pm f_{d}$. Finally, the global angular dependence of the DPDM signal will vary from one coherence time to the next, as the amplitudes $A^{\prime}_{m}$ stochastically vary. The global angular dependence of the axion signal, however, does not change as $a_{0}$ stochastically varies. Instead, the axion signal will slowly drift along with $\bm{B}_{\oplus}$. We note that $C_{10}$ only changes by $\mathcal{O}(1\%)$ over the 23-year duration of the high-fidelity dataset, so this is a much slower timescale.
@@ III SuperMAG high-fidelity dataset
The SuperMAG collaboration maintains a large database of three-axis magnetometer measurements from stations dispersed across the globe, which it makes publicly available for research purposes [54, 55, 56]. In this work, we search their “high-fidelity” dataset which reports magnetic-field measurements at 1-second intervals. This dataset contains many fewer stations than their “low-fidelity” (1-minute time resolution) dataset, which was analyzed in Refs. [50, 51], as not all stations report higher-frequency data. Likewise it spans a shorter period of time than the low-fidelity dataset. Many of the features of the high-fidelity dataset and the data-processing techniques applied to it are similar to those of the low-fidelity dataset, which were outlined in Ref. [50]. Here we will briefly review the important features, but we refer the interested reader to that prior work for a more detailed summary, or to Ref. [56] for a thorough description of the SuperMAG collaboration’s data-processing techniques.
@@
SuperMAG applies a rigorous data-cleaning procedure to the raw data it receives from its collaborators (prior to rotating coordinates). This procedure consists of three steps: automatic removal of data spikes; manual remove of errors; and automatic correction for instrument offsets/biases. First, spikes of duration five seconds or less are identified, removed, and replaced by linear interpolation. Next, manual identification and correction of other data errors are performed by highly experienced personnel. Lastly, instruments typically exhibit time-dependent offsets in each component that are corrected using known main-field models. Any temperature-dependent gain factors exhibited by the magnetometers are corrected for as well.
@@
Finally, we note that we have explicitly removed a few obvious anomalies from the dataset, specifically for the purpose of this analysis. The reported measurements in the dataset exceed $10^{5}\,\mathrm{nT}$ in six instances, in each case only for one or two seconds and immediately preceeding a period of inactivity. Due to the anomalously large size and short duration of these measurements, we believe them to be artifacts of the post-processing procedure. We replace all such measurements with null values (as if the station were inactive). In addition, the data reported by the single station located at Lucky Lake (IAGA code: LCL) exhibit a number of single-second spikes. These spikes have low amplitude, so that our analysis would actually upweight this station’s measurements (see discussion around Eqs. (19) and (20) for weighting procedure based on white noise level). This station is therefore unsuitable for common-noise modeling treatment with the other stations. Due to the prevalence and duration of these spikes, we do not suspect them to be physical magnetic-field measurements. In lieu of attempting to remedy these anomalies in post-processing, we simply discard the data from this single station.
@@ IV Analysis details
In this section, we describe the analysis of the high-fidelity dataset. This analysis is based on the analyses performed in Refs. [50, 51] on the low-fidelity dataset. The analyses in those works were complicated by the fact that the SuperMAG dataset consists of data from many stations which are active at different times. The strategy in Refs. [50, 51] was to first combine the time series of measurements from all $\mathcal{O}(500)$ stations into a handful of specific linear combinations whose forms were motivated by Eqs. (1) and (5). Refs. [50, 51] then partitioned these constructed time series into chunks of length roughly equal to the coherence time $T_{\mathrm{coh}}$ [see discussion around Eq. (2)] and searched for the DM signals in each coherence chunk independently. In order to perform this search, it was important to both understand how a potential signal would appear in these constructed time series and to characterize the noise levels of these time series. The former was done by applying the same combination procedure which was used on the magnetometer data to the signals in Eqs. (1) and (5), while the latter was estimated via a data-driven procedure. Once the expected signal size and noise levels were understood, Refs. [50, 51] computed a posterior distribution on the coupling $\varepsilon$ or $g_{a\gamma}$, which was then used to either set constraints on these parameters or identify signal candidates.
@@
We now describe the analysis procedure in detail, paying special attention to these new features which differ from Refs. [50, 51].
@@ IV.1 Time series
Each SuperMAG station $i$, located at coordinates $\Omega_{i}=(\theta_{i},\phi_{i})$, reports a time series of three-axis magnetic-field measurements $\bm{B}(\Omega_{i},t_{j})$ for some set of sampling times $t_{j}\in\mathcal{T}_{i}$. The set $\mathcal{T}_{i}$ of sampling times at which a particular station is active differs from one station to the next, making the prospect of analyzing stations separately and combining their results challenging. Instead, as in Refs. [50, 51], we first combine the time series of measurements from the individual stations into a handful of specific linear combinations, and then analyze these constructed time series as our input.
The manner of combining the time series will be motivated by the signals in Eqs. (1) and (5). That is, in the DPDM case, we weight the magnetic-field measurements from each station by the real and imaginary parts of the components of the $\bm{\Phi}_{1m}$ VSH
@@
On the other hand, in the axion case, we weight them by the components of the linear combination in Eq. (5)
@@
We can then define time series which combine the weighted measurements from different stations
@@
The sum here is over all stations $i$ which are active at time $t_{j}$. As in Refs. [50, 51], the station weights $w_{i}^{(n)}(t_{j})$ should be constant within a stationarity period (a week in the case of this work’s analysis), but may vary from one period to the next. In order to downweight noisier stations, we set the weights to be the inverse of the white noise level of the station; that is, for all $t$ within stationarity period $a$, we take
@@
Note that although the station weights $w_{i}^{(n)}$ are constant within a stationarity period, the total weights may not be because the number of active stations over which this sums run may vary within a stationarity period.
@@ IV.2 Data vector
Once the time series $X_{i}^{(n)}$ have been constructed, we will deal solely with these projected variables rather than the data from individual stations. Our goal is to search these time series for a nearly monochromatic signal. A natural strategy would therefore be to Fourier transform the data and look for excess power at a particular frequency. However, as described in Sec. II, the signal can only be treated as monochromatic within a coherence time $T_{\mathrm{coh}}\sim 10^{6}/f_{\mathrm{DM}}$. Therefore, we will instead partition the time series into “coherence chunks” each roughly of length $T_{\mathrm{coh}}$, and search each chunk independently for a monochromatic signal. We will then combine the results of these searches incoherently (i.e., we will treat them as independent searches).
@@
Because the coherence time $T_{\mathrm{coh}}$ depends explicitly on the DM frequency, this strategy should, in principle, require a different partition of the time series for each DM frequency we wish to search for. Such a procedure would be computationally intractable. Instead, Ref. [50] developed a procedure for choosing an approximate coherence time which could apply to a range of DM frequencies. In this way, the entire set of DM frequencies to be searched for could be split into $\mathcal{O}(50)$ ranges, each range could be assigned an approximate coherence time (which would differ from the true $T_{\mathrm{coh}}$ by at most 3% for all frequencies in the range), and the same partition of the time series could be used for the entire range. In this work, we apply the same procedure for selecting the DM frequencies to be searched for and determining the approximate coherence times. We restrict our analysis to $10^{-3}\,\textrm{Hz}\leq f_{\mathrm{DM}}\leq 0.98\,\textrm{Hz}$, [^8] We choose this upper bound because our analysis of the noise statistics in Appendix B.2 computes these statistics by averaging (in log-space) over a window of size $\sigma_{\log}=0.02$. We therefore cannot verify our assumptions on the noise statistics at frequencies $f_{\mathrm{DM}}\gtrsim 0.98\,\textrm{Hz}$. which requires $106$ frequency ranges. Rather than reiterating the details of this procedure, we refer the interested reader to Sec. V E of Ref. [50]. Henceforth, we will focus on the search for a single DM frequency $f_{\mathrm{DM}}$; we will simply refer to its approximate coherence time as $T_{\mathrm{coh}}$.
@@
Once we have fixed the coherence time $T_{\mathrm{coh}}$, we partition the full time series $X^{(n)}(t)$ into coherence chunks $X^{(n)}_{k}(t)$ each of length $T_{\mathrm{coh}}$. For each $k$, we wish to combine all the information in the time series $X^{(n)}_{k}$ which is relevant to searching for a DM signal of frequency $f_{\mathrm{DM}}$ into a single vector. Note that in the DPDM case, the signal may contain information at the frequency $f=f_{A^{\prime}}$, as well as $f=f_{A^{\prime}}\pm f_{d}$. Therefore, our data vector should include the Fourier transforms of all five time series at all three of these frequencies, that is we should define the 15-dimensional vector [^9] As $f_{d}$ may not generically be a discrete Fourier transform (DFT) frequency, it may be computationally difficult to compute these Fourier transforms at $f=f_{A^{\prime}}\pm f_{d}$. (The prescription for choosing frequencies outlined in Sec. V E of Ref. [50] is designed so that $f_{A^{\prime}}$ is always a DFT frequency.) As a result, we instead evaluate them at $f=f_{A^{\prime}}\pm\hat{f}_{d}$, where $\hat{f}_{d}$ is the DFT frequency closest to $f_{d}$., [^10] Throughout this section, $\vec{x}$ denotes a 15-dimensional vector in the DPDM case and a two-dimensional vector in the axion case. Meanwhile, $\bm{y}$ denotes a vector with three (typically spatial) dimensions.
@@
(See footnote 4 for the relationship between the Cartesian components $A^{\prime}_{i}$ and polarization amplitudes $A^{\prime}_{m}$.) These are normalized so that Eq. (2) becomes $\langle|\bm{c}_{k}|^{2}\rangle=1$. We refer the reader to Eqs. (36), (C1), and (C2) of Ref. [50] for the explicit expressions for $\vec{\mu}_{ik}$. In the axion case, the expectation value of $\vec{X}_{k}$ (in the presence of a signal) is

### v2

@@ I Introduction
Discerning the fundamental nature of dark matter (DM) remains one of the most prominent open problems in modern physics. A number of potential DM candidates have been proposed, but one class that has recently received significant attention is ultralight bosonic DM [1, 2]. This class includes DM candidates with masses $m_{\mathrm{DM}}\lesssim 1\,\textrm{eV}$, [^1] We work in natural units where $\hbar=c=1$. which have correspondingly large number densities and so behave as classical fields [3, 4, 5]. Two such candidates that have garnered particular interest are dark-photon dark matter (DPDM) [6, 7, 8, 9] and axionlike-particle dark matter (henceforth simply axion DM) [10, 11, 12, 13, 14]. Both of these candidates may exhibit couplings with electromagnetism [6, 15], [^2] In the context of grand unified theories (GUTs) or heterotic string theories, an axionlike particle that couples to photons should also possess a coupling to gluons, which would contribute to its mass [16, 17]. The discovery of a light axionlike particle coupled to photons would therefore rule out many simple GUT and string theory models. and a number of experimental [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36], astrophysical [37, 38, 39, 40, 41, 42, 43, 44, 45], and cosmological probes [46, 47, 48, 49, 50, 51] have been utilized to constrain their parameter spaces.
Few experiments searching for these ultralight DM candidates are sensitive to the very low mass parameter space $m_{\mathrm{DM}}\lesssim 10^{-12}\,\textrm{eV}$. This is because the signal in these experiments typically scales with the size of the experiment $L$, when $\lambda_{\mathrm{DM}}\gg L$ (where $\lambda_{\mathrm{DM}}=2\pi/m_{\mathrm{DM}}$ is the Compton wavelength of the DM). One notable example is a series of experimental searches that utilize Earth as a transducer for ultralight DM [52, 53, 54, 55, 56]. In such contexts, the effective size of the experiment is the radius of Earth $R$, which leads to a large enhancement of the signal.
@@
In this work, we extend the analyses of Refs. [53, 54] to the SuperMAG “high-fidelity” dataset, which exhibits a shorter one-second time resolution, albeit with fewer, $\mathcal{O}(200)$, total stations. This will enable sensitivity to Compton frequencies up to $f_{\mathrm{DM}}\leq 1\,\textrm{Hz}$. As ambient geomagnetic field noise is typically reduced at higher frequencies, this DM search strategy should be even more powerful at these higher frequencies. In fact, the SNIPE Hunt collaboration recently conducted an independent search for this same DM signal at even higher frequencies ($0.5\,\textrm{Hz}\leq f_{\mathrm{DM}}\leq 5\,\textrm{Hz}$) and was able to achieve competitive results with data from only three magnetically quiet locations [55, 56]. The search in this work leverages the benefits of both low noise at higher frequencies and the large amounts of data that the high-fidelity dataset has to offer. In this work, we search for ultralight DM in the frequency range $10^{-3}\,\textrm{Hz}\leq f_{\mathrm{DM}}\leq 0.98\,\textrm{Hz}$; which complements the existing searches from the SuperMAG low-fidelity dataset and the SNIPE Hunt collaboration.
This work is organized as follows. In Sec. II, we briefly review the magnetic-field signal of ultralight DM that we search for. In Sec. III, we introduce the SuperMAG high-fidelity dataset and review the data-processing techniques that have been applied to it. In Sec. IV, we describe the details of our analysis procedure. The analysis in this work largely follows the same procedure as in Refs. [53, 54] but with important modifications that we highlight in this section. In Sec. V, we present and discuss the resulting constraints on ultralight DM from our search. In Appendix A, we review the properties of vector spherical harmonics (VSH). In Appendix B, we validate our choice of analysis parameters by examining the statistics of the noise in our data. All the code used in this work is publicly available on Github [60].
@@ II Signal
Here $\Omega=(\theta,\phi)$ denotes the geographic coordinates on Earth’s surface, oriented so that $\theta=0$ corresponds to the Geographic North Pole and $\theta=\pi$ to the Geographic South Pole. (Note that while $\phi$ coincides with the usual definition of longitude, $\theta$ is not the usual latitude.) Additionally, $R$ is the radius of Earth, $A^{\prime}_{m}$ are complex amplitudes corresponding to the polarization modes of the dark photon, [^5] In our conventions, these amplitudes relate to the Cartesian components of $\bm{A}^{\prime}$ via $A^{\prime}_{x}=-\frac{1}{\sqrt{2}}\left(A^{\prime}_{+}-A^{\prime}_{-}\right),\quad A^{\prime}_{y}=-\frac{i}{\sqrt{2}}\left(A^{\prime}_{+}+A^{\prime}_{-}\right),\quad A^{\prime}_{z}=A^{\prime}_{0}.$ and $\bm{\Phi}_{\ell m}$ are VSH (see Appendix A). The presence of $f_{d}=1/(\textrm{sidereal day})$ in the time dependence of the signal arises due to the rotation of Earth. Here the dark-photon amplitudes $A^{\prime}_{m}$ are defined in the inertial celestial frame, while the signal in Eq. (1) is measured in the rotating Earth-fixed frame. Translating between these frames induces a polarization-dependent shift to the frequency of the signal. (In frequency space, this will introduce sidebands at $f=f_{A^{\prime}}\pm f_{d}$ to the signal.)
@@
Since Earth’s field slowly drifts over time, the Gauss coefficients $g_{\ell m}$ and $h_{\ell m}$ are time dependent. The IGRF-13 model provides values for these coefficients at five-year intervals from 1900 to 2020, along with their current time derivative. These values can be interpolated to find the values of the Gauss coefficients at any time between the beginning of 1900 and the beginning of 2020, and the time derivatives can be used to extrapolate their values up to the end of 2024.
@@
Before proceeding, we note a few distinctions between the signals in Eqs. (1) and (5). First, the DPDM signal is purely dipolar (only $\ell=1$ modes contribute). Because $\bm{B}_{\oplus}$ is approximately dipolar, the expansion in Eq. (3) is dominated by the $C_{10}$ term, but other terms contribute at the $\mathcal{O}(10\%)$ level. Likewise, the axion signal in Eq. (5) is predominantly dipolar, but receives small contributions from higher $\ell$ modes. Secondly, the axion signal in Eq. (5) does not feature any dependence on $f_{d}$. This is because the angular dependence of the axion signal is set by $\bm{B}_{\oplus}$, which corotates with Earth, as opposed to the DPDM polarization which does not. This means that, in frequency space, the axion signal will only appear at $f=f_{a}$, whereas the DPDM signal will appear at $f=f_{A^{\prime}},f_{A^{\prime}}\pm f_{d}$. Finally, the global angular dependence of the DPDM signal will vary from one coherence time to the next, as the amplitudes $A^{\prime}_{m}$ stochastically vary. The global angular dependence of the axion signal, however, does not change as $a_{0}$ stochastically varies. Instead, the axion signal will slowly drift along with $\bm{B}_{\oplus}$. We note that $C_{10}$ only changes by $\mathcal{O}(1\%)$ over the 23-year duration of the high-fidelity dataset, so this is a much slower timescale.
@@ III SuperMAG high-fidelity dataset
The SuperMAG collaboration maintains a large database of three-axis magnetometer measurements from stations dispersed across the globe, which it makes publicly available for research purposes [57, 58, 59]. In this work, we search their “high-fidelity” dataset which reports magnetic-field measurements at one-second intervals. This dataset contains many fewer stations than their “low-fidelity” (one-minute time resolution) dataset, which was analyzed in Refs. [53, 54], as not all stations report higher-frequency data. Likewise it spans a shorter period of time than the low-fidelity dataset. Many of the features of the high-fidelity dataset and the data-processing techniques applied to it are similar to those of the low-fidelity dataset, which were outlined in Ref. [53]. Here we will briefly review the important features, but we refer the interested reader to that prior work for a more detailed summary, or to Ref. [59] for a thorough description of the SuperMAG collaboration’s data-processing techniques.
@@
SuperMAG applies a rigorous data-cleaning procedure to the raw data it receives from its collaborators (prior to rotating coordinates). This procedure consists of three steps: automatic removal of data spikes; manual removal of errors; and automatic correction for instrument offsets/biases. First, spikes of duration five seconds or less are identified, removed, and replaced by linear interpolation. Next, manual identification and correction of other data errors are performed by highly experienced personnel. Last, instruments typically exhibit time-dependent offsets in each component that are corrected using known main-field models. Any temperature-dependent gain factors exhibited by the magnetometers are corrected for as well.
@@
Finally, we note that we have explicitly removed a few obvious anomalies from the dataset, specifically for the purpose of this analysis. The reported measurements in the dataset exceed $10^{5}\,\mathrm{nT}$ in six instances, in each case only for one or two seconds and immediately preceding a period of inactivity. Due to the anomalously large size and short duration of these measurements, we believe them to be artifacts of the post-processing procedure. We replace all such measurements with null values (as if the station were inactive). In addition, the data reported by the single station located at Lucky Lake (IAGA code: LCL) exhibit a number of single-second spikes. These spikes have low amplitude, so that our analysis would actually upweight this station’s measurements (see discussion around Eqs. (19) and (20) for weighting procedure based on white noise level). This station is therefore unsuitable for common-noise modeling treatment with the other stations. Due to the prevalence and duration of these spikes, we do not suspect them to be physical magnetic-field measurements. In lieu of attempting to remedy these anomalies in post-processing, we simply discard the data from this single station.
@@ IV Analysis details
In this section, we describe the analysis of the high-fidelity dataset. This analysis is based on the analyses performed in Refs. [53, 54] on the low-fidelity dataset. The analyses in those works were complicated by the fact that the SuperMAG dataset consists of data from many stations that are active at different times. The strategy in Refs. [53, 54] was to first combine the time series of measurements from all $\mathcal{O}(500)$ stations into a handful of specific linear combinations whose forms were motivated by Eqs. (1) and (5). Refs. [53, 54] then partitioned these constructed time series into chunks of length roughly equal to the coherence time $T_{\mathrm{coh}}$ [see discussion around Eq. (2)] and searched for the DM signals in each coherence chunk independently. In order to perform this search, it was important to both understand how a potential signal would appear in these constructed time series and to characterize the noise levels of these time series. The former was done by applying the same combination procedure that was used on the magnetometer data to the signals in Eqs. (1) and (5), while the latter was estimated via a data-driven procedure. Once the expected signal size and noise levels were understood, Refs. [53, 54] computed a posterior distribution on the coupling $\varepsilon$ or $g_{a\gamma}$, which was then used to either set constraints on these parameters or identify signal candidates.
@@
We now describe the analysis procedure in detail, paying special attention to these new features that differ from Refs. [53, 54].
@@ IV.1 Time series
Each SuperMAG station $i$, located at coordinates $\Omega_{i}=(\theta_{i},\phi_{i})$, reports a time series of three-axis magnetic-field measurements $\bm{B}(\Omega_{i},t_{j})$ for some set of sampling times $t_{j}\in\mathcal{T}_{i}$. The set $\mathcal{T}_{i}$ of sampling times at which a particular station is active differs from one station to the next, making the prospect of analyzing stations separately and combining their results challenging. Instead, as in Refs. [53, 54], we first combine the time series of measurements from the individual stations into a handful of specific linear combinations and then analyze these constructed time series as our input.
The manner of combining the time series will be motivated by the signals in Eqs. (1) and (5). That is, in the DPDM case, we project the magnetic-field measurements from each station onto the real and imaginary parts of the components of the $\bm{\Phi}_{1m}$ VSH
@@
On the other hand, in the axion case, we project them onto the components of the linear combination in Eq. (5)
@@
We can then define time series by a weighted combination of the projections from all the stations
@@
The sum here is over all stations $i$ that are active at time $t_{j}$. As in Refs. [53, 54], the station weights $w_{i}^{(n)}(t_{j})$ should be constant within a stationarity period (a week in the case of this work’s analysis) but may vary from one period to the next. In order to downweight noisier stations, we set the weights to be the inverse of the white noise level of the station; that is, for all $t$ within stationarity period $a$, we take
@@
Note that, although the station weights $w_{i}^{(n)}$ are constant within a stationarity period, the total weights may not be because the number of active stations over which this sum runs may vary within a stationarity period.
@@ IV.2 Data vector
Once the time series $X_{i}^{(n)}$ have been constructed, we will deal solely with these projected variables rather than the data from individual stations. Our goal is to search these time series for a nearly monochromatic signal. A natural strategy would therefore be to Fourier transform the data and look for excess power at a particular frequency. However, as described in Sec. II, the signal can only be treated as monochromatic within a coherence time $T_{\mathrm{coh}}\sim 10^{6}/f_{\mathrm{DM}}$. Therefore, we will instead partition the time series into “coherence chunks” each roughly of length $T_{\mathrm{coh}}$ and search each chunk independently for a monochromatic signal. We will then combine the results of these searches incoherently (i.e., we will treat them as independent searches).
@@
Because the coherence time $T_{\mathrm{coh}}$ depends explicitly on the DM frequency, this strategy should, in principle, require a different partition of the time series for each DM frequency we wish to search for. Such a procedure would be computationally intractable. Instead, Ref. [53] developed a procedure for choosing an approximate [^9] In Ref. [53] and in this work, this approximation of the coherence time holds to 3%. We note that the notion of a coherence time is itself approximate, as it is determined by $v_{\mathrm{DM}}$, which is known even less precisely than this. Therefore the use of this approximation (and the choice of a 3% tolerance) introduces no more error into our analysis than our use of the coherence time in the first place. coherence time that could apply to a range of DM frequencies. In this way, the entire set of DM frequencies to be searched for could be split into $\mathcal{O}(50)$ ranges, each range could be assigned an approximate coherence time (which would differ from the true $T_{\mathrm{coh}}$ by at most 3% for all frequencies in the range), and the same partition of the time series could be used for the entire range. In this work, we apply the same procedure for selecting the DM frequencies to be searched for and determining the approximate coherence times. We restrict our analysis to $10^{-3}\,\textrm{Hz}\leq f_{\mathrm{DM}}\leq 0.98\,\textrm{Hz}$, [^10] We choose this upper bound because our analysis of the noise statistics in Appendix B.2 computes these statistics by averaging (in log-space) over a window of size $\sigma_{\log}=0.02$. We therefore cannot verify our assumptions on the noise statistics at frequencies $f_{\mathrm{DM}}\gtrsim 0.98\,\textrm{Hz}$. which requires $106$ frequency ranges. Rather than reiterating the details of this procedure, we refer the interested reader to Sec. V E of Ref. [53]. Henceforth, we will focus on the search for a single DM frequency $f_{\mathrm{DM}}$; we will simply refer to its approximate coherence time as $T_{\mathrm{coh}}$.
@@
Once we have fixed the coherence time $T_{\mathrm{coh}}$, we partition the full time series $X^{(n)}(t)$ into coherence chunks $X^{(n)}_{k}(t)$ each of length $T_{\mathrm{coh}}$. For each $k$, we wish to combine all the information in the time series $X^{(n)}_{k}$ that is relevant to searching for a DM signal of frequency $f_{\mathrm{DM}}$ into a single vector. Note that in the DPDM case, the signal may contain information at the frequency $f=f_{A^{\prime}}$, as well as $f=f_{A^{\prime}}\pm f_{d}$. Therefore, our data vector should include the Fourier transforms of all five time series at all three of these frequencies, that is we should define the 15-dimensional vector [^11] As $f_{d}$ may not generically be a discrete Fourier transform (DFT) frequency, it may be computationally difficult to compute these Fourier transforms at $f=f_{A^{\prime}}\pm f_{d}$. (The prescription for choosing frequencies outlined in Sec. V E of Ref. [53] is designed so that $f_{A^{\prime}}$ is always a DFT frequency.) As a result, we instead evaluate them at $f=f_{A^{\prime}}\pm\hat{f}_{d}$, where $\hat{f}_{d}$ is the DFT frequency closest to $f_{d}$., [^12] Throughout this section, $\vec{x}$ denotes a 15-dimensional vector in the DPDM case and a two-dimensional vector in the axion case. Meanwhile, $\bm{y}$ denotes a vector with three (typically spatial) dimensions.
@@
(See footnote 5 for the relationship between the Cartesian components $A^{\prime}_{i}$ and polarization amplitudes $A^{\prime}_{m}$.) These are normalized so that Eq. (2) becomes $\langle|\bm{c}_{k}|^{2}\rangle=1$. We refer the reader to Eqs. (36), (C1), and (C2) of Ref. [53] for the explicit expressions for $\vec{\mu}_{ik}$. In the axion case, the expectation value of $\vec{X}_{k}$ (in the presence of a signal) is